OpenStudy (anonymous):

for which equation is x=9 a solution? a. 3x-7=-34 b. x-9=1 c. x/9-8=9 d. 4x+5=41

OpenStudy (anonymous):

B

OpenStudy (diyadiya):

D

OpenStudy (anonymous):

how is it D

OpenStudy (diyadiya):

4*9 + 5= 36 + 5 = 41